OpenStudy (anonymous):

please help me x^2-___+49/4=( )^2

OpenStudy (anonymous):

so what is supposed to be where the ___ is?

OpenStudy (anonymous):

that is how the question is asked it is a 2 part answer x^2-what+49/9=( )^2

OpenStudy (anonymous):

49/4

OpenStudy (mertsj):

7x

OpenStudy (mertsj):

\[x^2-7x+\frac{49}{4}=(x-\frac{7}{2})^2\]

OpenStudy (anonymous):